### v1

# TorchTitan: One-stop PyTorch native solution for production ready LLM pre-training

Abstract
The development of large language models (LLMs) has been instrumental in advancing state-of-the-art natural language processing applications. Training LLMs with billions of parameters and trillions of tokens require sophisticated distributed systems that enable composing and comparing several state-of-the-art techniques in order to efficiently scale across thousands of accelerators. However, existing solutions are complex, scattered across multiple libraries/repositories, lack interoperability, and are cumbersome to maintain. Thus, curating and empirically comparing training recipes require non-trivial engineering effort.
This paper introduces TorchTitan, an open-source, PyTorch-native distributed training system that unifies and advances state-of-the-art techniques, streamlining integration and reducing engineering overhead. TorchTitan enables seamless application of 3D parallelism in a modular and composable manner, while featuring elastic scaling to adapt to changing computational requirements. The system provides comprehensive logging, efficient checkpointing, and debugging tools, ensuring production-ready training. Moreover, TorchTitan incorporates innovative hardware-software co-designed solutions, leveraging cutting-edge features like Float8 training and SymmetricMemory to maximize hardware utilization. As a flexible experimental test bed, TorchTitan facilitates the curation and comparison of custom recipes for diverse training contexts. By leveraging TorchTitan, we developed optimized training recipes for the Llama 3.1 family and provide actionable guidance on selecting and combining distributed training techniques to maximize training efficiency, based on our hands-on experiences.
We thoroughly assess TorchTitan on the Llama 3.1 family of LLMs, spanning 8 billion to 405 billion parameters, and showcase its exceptional performance, modular composability, and elastic scalability. By stacking training optimizations, we demonstrate accelerations of 65.08% with 1D parallelism at the 128-GPU scale (Llama 3.1 8B), an additional 12.59% with 2D parallelism at the 256-GPU scale (Llama 3.1 70B), and an additional 30% with 3D parallelism at the 512-GPU scale (Llama 3.1 405B) on NVIDIA H100 GPUs over optimized baselines.
[…]
1 Introduction
LLMs are at the forefront of NLP advancement. Large Language Models (LLMs) (Devlin, 2018; Liu et al., 2019; Radford et al., 2019; Chowdhery et al., 2023; Anil et al., 2023; Achiam et al., 2023; Dubey et al., 2024; Jiang et al., 2024; Abdin et al., 2024) have been driving force behind the advancement of natural language processing (NLP) applications spanning language translation, content/code generation, conversational AI, text data analysis, creative writing and art, education and research etc.
LLMs require billions of parameters and training over trillion tokens to achieve state-of-the-art performance. Achieving state-of-the-art LLM performance requires massive scale, exemplified by top-performing models like Llama 3.1 (405B parameters, 15T tokens, 30.84M GPU hours, 16K H100 GPUs) (Dubey et al., 2024) and Google’s PaLM (540B parameters, 0.8T tokens, 9.4M TPU hours, 6144 TPUv4 chips) (Chowdhery et al., 2023). These models demonstrate exceptional natural language understanding and generation capabilities, but necessitate substantial computational resources, memory, and time to train, highlighting the significant investment required to advance natural language processing.
LLM training challenges are being tackled from all sides. Training large language models (LLMs) at scale is a daunting task that requires a delicate balance of parallelism, computation, and communication, all while navigating intricate memory and computation tradeoffs. The massive resources required for training make it prone to GPU failures, underscoring the need for efficient recovery mechanisms and checkpointing strategies to minimize downtime (Eisenman et al., 2022; Wang et al., 2023; Gupta et al., 2024; Maurya et al., 2024; Wan et al., 2024). To optimize resource utilization and achieve elastic scalability, it is crucial to combine multiple parallelism techniques, including Data Parallel (Li et al., 2020; Rajbhandari et al., 2020; Zhang et al., 2022; Zhao et al., 2023), Tensor Parallel (Narayanan et al., 2021; Wang et al., 2022; Korthikanti et al., 2023), Context Parallel (Liu et al., 2023; Liu and Abbeel, 2024; NVIDIA, 2023; Fang and Zhao, 2024), and Pipeline Parallel (Huang et al., 2019; Narayanan et al., 2019, 2021; Tang et al., 2024). By stacking these parallelisms with memory and computation optimization techniques, such as activation recomputation (Chen et al., 2016; Korthikanti et al., 2023; He and Yu, 2023), mixed precision training (Micikevicius et al., 2018, 2022), and deep learning compilers (Bradbury et al., 2018; Yu et al., 2023; Li et al., 2024; Ansel et al., 2024), it is possible to maximize hardware utilization.
Limitations of existing systems incorporating state-of-the-art techniques. While state-of-the-art distributed training techniques have significantly advanced the field, existing systems that incorporate them still fall short in addressing critical challenges that hinder their usability, adoption and effectiveness for researchers and industry practitioners.
  1. 1.
    Non-composable: Existing systems struggle to combine and stack various parallelism techniques, limiting the exploration of multi-dimensional parallelism. Further integrating them with memory and computation optimizations is challenging, hindering training efficiency.
  2. 2.
    Inflexible and monolithic architecture: Current systems are not modular or extensible, making it difficult to integrate and compare new techniques, optimizations, and hardware, and limiting adaptability to evolving machine learning landscapes.
  3. 3.
    Inefficient hardware utilization: Current systems fail to fully leverage advanced hardware features, leading to sub-optimal GPU efficiency, and lack customizable activation checkpointing strategies to navigate memory-computation trade-offs.
  4. 4.
    Insufficient support for production-grade training: Existing systems lack scalable and efficient distributed checkpointing, making failure recovery and model saving cumbersome, and often do not provide adequate debugging tools and logging metrics, leading to difficulties in identifying and fixing issues, particularly for those without extensive expertise.
  5. 5.
    Existing systems fall short in harnessing the full potential of frameworks like PyTorch, missing out on bug fixes, optimized kernels, new features, and compiler support. They also rely on external dependencies that often lack thorough testing and can become outdated or incompatible due to inadequate maintenance.
Root cause: Lack of an expressive tensor abstraction. The root cause of non-composability and inflexibility of a distributed system stems from the lack of using an expressive tensor and device abstraction as a central component, upon which all of the distributed parallelisms, checkpointing, and efficiency optimizations can be built.
Design Principle: Unified distributed tensor and device abstractions as building blocks. A unified device abstraction represents the distributed system as a multi-dimensional array, where each dimension corresponds to a parallelism technique, managing communication between devices and handling collective process groups. A complementary tensor abstraction enables tensors to be sharded across this array, maintaining sharding specifications and supporting automatic sharding propagation. Together, these abstractions enable seamless composition of parallelism techniques, ensure correct semantics, and facilitate dispatching of collectives for distributed operations
We address the technical challenge of a unified tensor abstraction by employing PyTorch’s Distributed Tensor (DTensor) and DeviceMesh (Wanchao Liang, 2023) as the foundational components for TorchTitan. Through our work with DTensor and DeviceMesh, we identified key limitations and addressed them. By using and extending DTensor, we develop TorchTitan, a production-ready system that enables composability, modularity, flexibility, and extensibility in distributed training. TorchTitan facilitates the composition of 3D parallelism, training optimizations, scalable distributed checkpointing, and harnesses the full benefits of the PyTorch ecosystem.
To develop and evaluate the capabilities of TorchTitan, we undertook several key steps, which represent the core contributions of this work, and are summarized as follows:
  1. 1.
    We advance DTensor by extending its sharding to support n-D parallelism, adding compatibility with torch.compile for compiler optimizations, and enabling efficient checkpointing of n-D models via state dict support. We also resolve critical bugs to bolster DTensor’s production readiness.
  2. 2.
    We demonstrate how to compose and stack various parallelism techniques, facilitating the exploration of multi-dimensional parallelism in large language model training (§2.1).
  3. 3.
    We enable novel hardware-software co-designed solutions exploiting advanced hardware features to increase GPU efficiency, offer customizable activation checkpointing strategies for navigating memory-computation trade-offs, and utilize torch.compile to further optimize memory, computation, and communication (§2.2).
  4. 4.
    We offer production grade training by incorporating scalable and efficient distributed checkpointing to facilitate fast failure recovery, integrating debugging tools like Flight Recorder to debug crashed/stuck jobs, and providing extensive logging metrics (§2.3).
  5. 5.
    We extensively evaluate TorchTitan on the Llama 3.1 family of models (8B, 70B, and 405B with 1D, 2D, and 3D parallelisms, respectively) at the scale from 8 to 512 GPUs to demonstrate elastic scalability while ensuring efficiency, convergence, and accuracy. In summary, we demonstrate training accelerations of 65.08% with 1D parallelism at the 128-GPU scale (Llama 3.1 8B), an additional 12.59% with 2D parallelism at the 256-GPU scale (Llama 3.1 70B), and an additional 30% with 3D parallelism at the 512-GPU scale (Llama 3.1 405B) on latest NVIDIA H100 GPUs over optimized baselines (§3.2).
  6. 6.
    We provide systematic training recipes and guidelines that empower users to navigate the complexities of distributed training, helping them optimize training efficiency for a range of model sizes and cluster configurations (§3.3).
  7. 7.
    We show how our modular and extensible architecture allows for seamless integration and comparison of new techniques, optimizations, and hardware, ensuring adaptability to evolving machine learning landscapes (§4).
By providing an accessible and extensible platform, TorchTitan democratizes large language model (LLM) pre-training, empowering a wider range of researchers and developers to tap into the potential of LLMs and accelerate innovation in the field.
2 Elasticity through Composability
Figure 1: Composable and Modular TorchTitan initialization workflow.
TorchTitan incorporates various parallelism techniques in a modular manner to enable easy, user-selectable combinations of multi-dimensional parallelisms. This composability enables the tackling of difficult scaling challenges by enhancing the ease of frontier exploration for optimizing training efficiencies at scale.
The codebase of TorchTitan is organized purposefully to enable composability and extensibility. We intentionally keep three main components separate and as orthogonal as possible: (1) the model definition, which is parallelism-agnostic and designed for readability, (2) parallelism helpers, which apply Data Parallel, Tensor Parallel, and Pipeline Parallel to a particular model, and (3) a generalized training loop. All these components are configurable via TOML files with command-line overrides, and it is easy to add new models and parallelism techniques on top of the existing codebase.
2.1 Composable N-D parallelism training
In this section, we will walk through the entire regime of scaling model training on large clusters, including meta device initialization and the core composable multi-dimensional parallelisms, to showcase how these techniques can be composed to train LLMs efficiently at increasing scale in TorchTitan. The corresponding actual code snippets in TorchTitan can be found in Appendix 8.
2.1.1 Large-scale model initialization using meta device
Given the exponential increase in model sizes for LLMs, the first scaling issue appears even before the actual training starts. This is the need to instantiate a large model for sharding across the cluster, yet without overflowing CPU or GPU memory.
To tackle this, we enabled meta device initialization for models in TorchTitan, where the model is first initialized on a “meta” device type. The meta device tensor only holds the metadata information, not the actual data, making initialization ultra-fast. After that, we perform model sharding and transforming the model parameters into Distributed Tensors (DTensors) where each parameter holds a local shard that lives on the meta device. Finally, we perform parameter initialization based on the user-defined initialization functions. We leverage Distributed Tensor to properly sync Random Number Generator (RNG) seeds, and initialize the parameters according to their sharding layouts. This ensures the parameters start with the same values as if the whole model were initialized on one device before sharding, and thus facilitating convergence comparisons between different parallelism configurations.
2.1.2 Fully Sharded Data Parallel
The original Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023) is an effective implementation of ZeRO that offers large model training capability in PyTorch. However, the original implementation (FSDP1) in PyTorch suffers from various limitations due to its FlatParameter implementation (see details in Appendix 9.1).
Given these limitations, TorchTitan integrates a new version of Fully Sharded Data Parallel (FSDP2), which uses the per-parameter Distributed Tensor sharding representation and thus provides better composability with model parallelism techniques and other features that require the manipulation of individual parameters,
TorchTitan integrates and leverages FSDP2 as it’s default 1D parallelism, benefiting from the improved memory management (often 7 percent lower per GPU memory requirement vs FSDP1) and the slight performance gains (average of 1.5 percent gain vs FSDP1). More details on FSDP2 and usage example are shown in Appendix 9.1. TorchTitan makes it simple to run with FSDP2 by embedding appropriate defaults, including auto-sharding with your world size automatically.
For scaling to even larger world sizes, TorchTitan also integrates Hybrid Sharded Data Parallel (HSDP) which extends FSDP2 by creating sharding groups. Details are shown in Appendix 9.2
2.1.3 Tensor Parallel
Tensor Parallel (TP) (Narayanan et al., 2021), together with Sequence Parallel (SP) (Korthikanti et al., 2023), is a key model parallelism technique to enable large model training at scale.
TP is implemented in TorchTitan using the PyTorch’s RowwiseParallel and ColwiseParallel APIs, where the model parameters are partitioned to DTensors and perform sharded computation with it. By leveraging DTensor, the TP implementation does not need to touch the model code, which allows faster enablement on different models and provides better composability with other features mentioned in this paper.
Tensor and Sequence Parallel (TP/SP)
While TP partitions the most computationally demanding aspects, Sequence Parallel (SP) perform sharded computation for the normalization or dropout layers on the sequence dimension, which otherwise generate large replicated activation tensors, and thus can be challenging to memory constraints per GPU. See Appendix 9.3 for more details, illustrations, and usage for both TP and FSDP + TP.
Because of the synergistic relationship between TP and SP, TorchTitan natively bundles these two together and they are jointly controlled by the TP degree setting.
Loss Parallel
When the loss function is computed, the model outputs are usually very large. Since the model outputs from TP/SP are sharded on the (often huge) vocabulary dimension, naively computing the cross-entropy loss requires gathering all the shards along the TP dimension to make the outputs be replicated, which incurs large memory usage.
With Loss Parallel, the cross entropy loss can be computed efficiently, without gathering all the model output shards to every single GPU. This not only significantly reduces the memory consumption, but also improves training speed by reducing communication overhead and doing sharded computation in parallel. Given these improvements, TorchTitan implements loss parallel by default.
2.1.4 Pipeline Parallel
For pre-training at the largest scales, TorchTitan offers Pipeline Parallelism, which becomes essential due to having the lightest amount of communication overhead and leveraging P2P communications.
Pipeline Parallel (PP) views the model as a sequence of operations, chunking the operations (and the parameters used by them) into S stages which run on separate groups of devices. In the typical case, one stage represents a single model layer or a group of N adjacent model layers, but in theory it could be even be a partial layer. For the forward pass, a stage receives input activations (except stage 0), performs local computation, and sends output activations (except stage S - 1). The last stage performs a loss computation, and begins the backward pass, sending gradients in the reverse order through the pipeline. To improve efficiency, the input batch is broken into microbatches and a pipeline schedule overlaps computation on one microbatch with communication for others. TorchTitan enables a variety of pipeline schedules, with their schedules previously described in other works (Narayanan et al., 2019; Huang et al., 2019; Narayanan et al., 2021; Tang et al., 2024).
The training loop must also account for creation of pipeline stages, and executing a pipeline schedule rather than invoking model.forward() directly. Because the schedule computes loss per microbatch, the loss computation and any logging code must be updated for PP. In TorchTitan, we propose to define a shared loss_fn to be used by both pipeline and non-pipeline code paths, and thus minimise divergence in the training loop.
Interactions with data parallelism, such as ensuring that data-parallel reduction happens only after the last microbatch in the schedule, and scheduling of shard and unshard operations when using zero-3, are also handled transparently inside the pipeline schedule executor, simplifying the trainer implementation in TorchTitan. For it’s usage in TorchTitan, please see Appendix 9.4.
2.2 Optimizing training efficiencies
2.2.1 Navigating compute-memory trade-offs using activation checkpointing
Activation checkpointing (AC) (Chen et al., 2016) and selective activation checkpointing (SAC) (Korthikanti et al., 2023) are standard training techniques to reduce peak GPU memory usage, by trading activation recomputation during the backward pass for memory savings. It is often needed even after applying multi-dimensional parallelisms.
TorchTitan offers flexible AC and SAC options utilizing torch.utils.checkpoint, applied at the TransformerBlock level. The AC strategies include “full” AC, op-level SAC, and layer-level SAC.
Within a TransformerBlock, full AC works by recomputing all activation tensors needed during the backward pass, whereas op-level SAC saves the results from computation-intensive PyTorch operations and only recomputes others. Layer-level SAC works in similar fashion as full AC, but the wrapping is applied to every $x$ TransformerBlock (where $x$ is specified by the user) to implement configurable trade-offs between memory and recompute. (Details are in Appendix  9.5.)
2.2.2 Regional compilation to exploit torch.compile optimizations
torch.compile was released in PyTorch 2 (Ansel et al., 2024) with TorchDynamo as the frontend to extract PyTorch operations into an FX graph, and TorchInductor as the backend to compile the FX graph into fused Triton code to improve the performance.
In TorchTitan, we use regional compilation, which applies torch.compile to each individual TransformerBlock in the Transformer model. This has two main benefits: (1) we get a full graph (without graph breaks) for each region, compatible with FSDP2 and TP (and more generally torch.Tensor subclasses such as DTensor) and other PyTorch distributed training techniques; (2) since the Llama model stacks identical TransformerBlock layers one after another, torch.compile can identify the same structure being repeatedly compiled and only compile once, thus greatly reducing compilation time.
torch.compile brings efficiency in both throughput and memory (see Section 3.2) via computation fusions and computation-communication reordering, in a model-agnostic way with a simple user interface. Below we further elaborate how torch.compile composability helps TorchTitan unlock hardware-optimized performance gain with simple user interface, with the integration of advanced features such as Asynchronous TP and Float8.
2.2.3 Asynchronous Tensor Parallel to maximally overlap communication
By default TP incurs blocking communications before/after the sharded computations, causing computation resources to not be effectively utilized. Asynchronous TP (AsyncTP) (Wang et al., 2022) achieves computation-communication overlap by fractionalizing the TP matrix multiplications within the attention and feed-forward modules into smaller chunks, and overlapping communication collectives in between each section. The overlap is achieved by a micro-pipelining optimization, where results are being communicated at the same time that the other chunks of the matmul are being computed.
PyTorch AsyncTP is based on a SymmetricMemory abstraction, which creates intra-node buffers to write faster communication collectives (Wang et al., 2024). This is done by allocating an shared memory buffer on each GPU in order to provide direct P2P access.
With TorchTitan’s integration of torch.compile, AsyncTP can be easily configured in TorchTitan to achieve meaningful end-to-end speedups (see Section 3.2 for details) on newer hardware (H100 or newer GPUs with NVSwitch within a node). Usage details are in Appendix 9.6
2.2.4 Boosting throughput with mixed precision training and Float8 support
Mixed precision training (Micikevicius et al., 2018) provides both memory and computational savings while ensuring training stability. FSDP2 has built-in support for mixed precision training with basic torch.dtype. This covers the popular usage of performing FSDP all-gather and computation in a low precision (e.g. torch.bfloat16), and perform lossless FSDP reduce-scatter (gradient) in high precision (e.g. torch.float32) for better numerical results. See Appendix  9.7 for usage details.
TorchTitan also supports more advanced mixed precision training with Float8 (a derived data type) on newer hardware like H100, with substantial performance gains (reported in Section 3.2). The Float8 feature from torchao.float8 supports multiple per-tensor scaling strategies, including dynamic, delayed, and static (see Micikevicius et al. (2022); Vasiliy Kuznetsov (2024), Section 4.3 for details), while being composable with other key PyTorch-native systems such as autograd, torch.compile, FSDP2 and TP (with Float8 all-gather capability (Feng et al., 2024)).
2.3 Production ready training
To enable production-grade training, TorchTitan offers seamless integration with key features out of the box. These include (1) efficient checkpointing using PyTorch Distributed Checkpointing (DCP), and (2) debugging stuck or crashed jobs through integration with Flight Recorder.
2.3.1 Scalable and efficient Distributed Checkpointing
Checkpoints are crucial in training large language models for two reasons: they facilitate model reuse in applications like inference and evaluation, and they provide a recovery mechanism in case of failures. An optimal checkpointing workflow should ensure ease of reuse across different parallelisms and maintain high performance without slowing down training. There are two typical checkpointing methods. The first aggregates the state (model parameters and optimizer states) into an unsharded version that is parallelism-agnostic, facilitating easy reuse but requiring expensive communication. The second method has each trainer save its local sharded state, which speeds up the process but complicates reuse due to embedded parallelism information.
DCP addresses these challenges using DTensor, which encapsulates both global and local tensor information independently of parallelism. DCP converts this information into an internal format for storage. During loading, DCP matches the stored shards with the current DTensor-based model parameters and optimizer states, fetching the necessary shard from storage. TorchTitan, which utilizes all native PyTorch parallelisms, effectively uses DCP to balance efficiency and usability. Furthermore, DCP enhances efficiency through asynchronous checkpointing by processing storage persistence in a separate thread, allowing this operation to overlap with subsequent training iterations. TorchTitan utilizes DCP’s asynchronous checkpointing to reduce the checkpointing overhead by 5-15x compared to synchronous distributed checkpointing for the Llama 3.1 8B model (Zhang et al., 2024; Huang et al., 2024).
2.3.2 Flight Recorder to debug job crashes
A common failure mode when developing parallelism code, or when running at large scale, is to observe a NCCL collective timeout and then the need to figure out the root cause. Since communication kernels are typically asynchronous from the perspective of the CPU, by the time something times out, it can be very hard to pinpoint which operation failed and why. PyTorch provides a Flight Recorder for NCCL collectives to help resolve this dilemma. It records the start and end time (on GPU) as well as the enqueue time (on CPU) for every collective or p2p operation. Additionally, it logs metadata such as which process group was used, who the source rank is (and destination, for p2p), tensor sizes, and stack traces.
We find the data contained in the Flight Recorder helps debug collective hangs and p2p hangs caused by bugs in parallelism code. For PP, there may be schedule bugs that lead to hangs, due to a missing or improperly ordered send or recv operation. Analysis based on the Flight Recorder data can pinpoint the latest send or recv that has completed on the GPU. For FSDP or TP, it is possible to determine whether one or more ranks has not called into a collective, perhaps due to a bug in PP scheduling or faulty logic in TP.
3 Experimentation
In this section, we demonstrate the effectiveness of elastic distributed training using TorchTitan, via experiments on Llama 3.1 8B, 70B, and 405B, from 1D parallelism to 3D parallelism (respectively), at the scale from 8 GPUs to 512 GPUs. We also share the knowledge and experience gained through TorchTitan experimentation. A walkthrough of the codebase on how we apply (up to) 3D parallelism can be found in Appendix 8.
3.1 Experimental setup
The experiments are conducted on NVIDIA H100 GPUs [^1] The H100 GPUs used for the experiments are non-standard. They have HBM2e and are limited to a lower TDP. The actual peak TFLOPs should be between SXM and NVL, and we don’t know the exact value. with 95 GiB memory, where each host is equipped with 8 GPUs and NVSwitch. Two hosts form a rack connected to a TOR switch. A backend RDMA network connects the TOR switches. In TorchTitan we integrate a checkpointable data loader and provide built-in support for the C4 dataset (en variant), a colossal, cleaned version of Common Crawl’s web crawl corpus (Raffel et al., 2020). We use the same dataset for all experiments in this section. For the tokenizer, we use the official one (tiktoken) released together with Llama 3.1.
3.2 Performance
To showcase the elasticity and scalability of TorchTitan, we experiment on a wide range of GPU scales (from 8 to 512), as the underlying model size increases (8B, 70B, and 405B) with a varying number of parallelism dimensions (1D, 2D, and 3D, respectively). To demonstrate the effectiveness of the optimization techniques introduced in Section 2.2, we show how training throughput improves when adding each individual technique on appropriate baselines. In particular, when training on a higher dimensional parallelism with new features, the baseline is always updated to include all previous techniques.
We note that, throughout our experimentation, memory readings are stable across the whole training process [^2] Different PP ranks can have different peak memory usages. We take the maximum across all GPUs., whereas throughput numbers (token per second, per GPU) are calculated and logged every 10 iterations, and always read at the (arbitrarily determined) 90th iteration. We do not report Model FLOPS Utilization (MFU) (Chowdhery et al., 2023) because when Float8 is enabled in TorchTitan, both BFLOAT16 Tensor Core and FP8 Tensor Core are involved in model training, but they have different peak FLOPS and the definition of MFU under such scenario is not well-defined. We note that the 1D Llama 3.1 8B model training on 8 or 128 H100 GPUs without Float8 achieves 33% to 42% MFU.
Table 1: 1D parallelism (FSDP) on Llama 3.1 8B model, 8 GPUs. Mixed precision training. Selective activation checkpointing. Local batch size 2, global batch size 16.
Techniques Throughput (Tok/Sec) Comparison Memory (GiB)
FSDP 6,258 100% 81.9
+ torch.compile 6,674 + 6.64% 77.0
+ torch.compile + Float8 9,409 + 50.35% 76.8
Table 2: 1D parallelism (FSDP) on Llama 3.1 8B model, 128 GPUs. Mixed precision training. Selective activation checkpointing. Local batch size 2, global batch size 256.
Techniques Throughput (Tok/Sec) Comparison Memory (GiB)
FSDP 5,645 100% 67.0
+ torch.compile 7,168 + 26.97% 62.0
+ torch.compile + Float8 9,319 + 65.08% 61.8
Table 3: 2D parallelism (FSDP + TP) + torch.compile + Float8 on Llama 3.1 70B model, 256 GPUs. Mixed precision training. Full activation checkpointing. FSDP degree 32, TP degree 8. Local batch size 16, global batch size 512.
Techniques Throughput (Tok/Sec) Comparison Memory (GiB)
2D 897 100% 70.3
+ AsyncTP 1,010 + 12.59% 67.7
Table 4: 3D parallelism (FSDP + TP + PP) + torch.compile + Float8 + AsyncTP on Llama 3.1 405B model, 512 GPUs. Mixed precision training. Full activation checkpointing. FSDP degree 4, TP degree 8, PP degree 16. Local batch size 32, global batch size 128.
Schedule Throughput (Tok/Sec) Comparison Memory (GiB)
1F1B 100 100% 78.0
Interleaved 1F1B 130 + 30.00% 80.3
3.3 Scaling with TorchTitan 3D parallelism
The LLM scaling law imposes challenges due to increasingly larger model size and huge amount of data, which requires applying parallelism strategies on massive number of GPUs. TorchTitan provides the ability to compose different parallelisms to efficiently scale model training to thousands of GPUs. This section discusses the observations and motivations to apply TorchTitan 3D parallelism when training LLMs at large scale. Please note that there could be many 3D parallelism combinations, but we choose to only discuss one combination in this paper, which could be summarized as the following diagram:
Figure 2: Scaling with 3D Parallelism
3.3.1 Scaling with FSDP
FSDP (ZeRO) is a generalized technique which can be applied to any model architecture, making it a good choice for the first or the only degree of parallelism. As long as the FSDP communication is faster than the corresponding computation (which is the case for LLMs trained on up to hundreds of, say 512, GPUs), and there is no need to reduce (effective) per-GPU batch size to be below 1 (for reasons mentioned below in the TP section), 1D FSDP should be sufficient.
Existing ring-based implementations of NCCL collectives (all-gather, reduce-scatter) will incur a latency overhead which becomes severe at large scale (e.g. 512 GPUs). FSDP alone will become less efficient due to the collective latency increasing linearly with the world size, resulting in FSDP collectives that cannot be hidden by the computation any more. To further scale out, one needs to consider combining model parallelism solutions such as TP and PP.
3.3.2 2D parallelism: Apply TP with FSDP
Model Parallelism (TP and PP) can help avoid the increased collective latency faced by scaling FSDP alone. TP can further lower the effective local batch size (to a minimum of 1TP $\frac{1}{\text{TP degree}}$ when the local batch size is set to 1), as TP sharded models on multiple GPUs work jointly to process the same batch. This can be vital for reducing peak memory usage so that training could fit in GPU memory (e.g. due to large model size or sequence length), or for strong scaling with fixed desired global batch size (e.g. due to training efficiency considerations).
In addition, TP performs feature dimension sharding. This can bring more optimized matrix multiplication shapes for better FLOP utilization.
As TP introduces extra blocking collectives, in practice TP is only applied within a node that have fast interconnect (NVLink). AsyncTP could improve the performance by fully overlapping the communication, but could not scale to multi-node easily, so the TP degree is usually limited to 8. This means to scale beyond an ultra-large number of GPUs (e.g. more than 4192 GPUs), we need 3D parallelism that combines PP.
3.3.3 3D Parallelism: Apply PP with 2D Parallelism
Compared to other model parallelisms, PP requires less communication bandwidth by virtue of only transmitting activations and gradients between stages in a P2P manner. It is especially useful (1) to further reduce FSDP communication latency when the FSDP world size becomes large again that FSDP+TP still exposes FSDP collectives; or (2) to train with bandwidth-limited cluster.
We note that, the performance of PP, and in particular the “bubble” size, could vary by pipeline schedules being used and the microbatch size, assuming fixed global batch size and the world size.
4 Demonstrating adaptability and extensibility
In this section, we demonstrate the adaptability and extensibility of TorchTitan by highlighting ongoing work and external contributions that showcase its ability to seamlessly integrate and compare new techniques, optimizations, and models.
4.1 Ongoing work: 4D parallelism and zero-bubble pipeline schedules
TorchTitan’s modular and extensible architecture enables the seamless integration of new techniques and optimizations. For instance, ongoing work includes incorporating Context Parallel (Liu et al., 2023; Liu and Abbeel, 2024; NVIDIA, 2023) to enable 4D parallelism, and leveraging the torch.distributed.pipelining package to support zero-bubble schedules (Tang et al., 2024). This demonstrates TorchTitan’s ability to adapt to evolving machine learning landscapes.
4.2 External contributions: Building and evaluating custom innovations
TorchTitan’s flexible architecture also empowers users to easily integrate and compare new innovations. By providing a modular and efficient test bed, TorchTitan enables users to rapidly benchmark new techniques, optimizations, and hardware on their training performance. This has already led to the refinement of a new production-grade dataloader, improvements in a new ZeRO implementation, advancements in an Adam-based optimizer, and the training of a top-tier diffusion model.
5 Related works
With the rapidly growing significance of LLMs  (Dubey et al., 2024; Achiam et al., 2023), there is substantial research and industry focus on improving infrastructure for training LLMs of various sizes. Since the very nature of these models are large, distributed training support becomes inevitable. Libraries like Megatron (Narayanan et al., 2021), DeepSpeed (Rasley et al., 2020), and PyTorch distributed (Pytorch native) (Paszke et al., 2019; Meta Platforms, Inc., 2024a) offer APIs to build distributed training workflows. NVIDIA NeMo (NVIDIA Corporation, 2024), built on Megatron-LM, offers a packaged solution for handling complex end-to-end model life-cycle from data curation to model deployment. Pytorch-native solutions like torchtune (Meta Platforms, Inc., 2024b) focus on fine-tuning LLMs in a simplified workflow. TorchTitan differs from these solutions by focusing on production grade pre-training using PyTorch-native APIs. The library is designed with elastic composability to accomodate the scale required to pre-train LLMs with minimal external dependencies. This lowers the bar to understand and extend pre-training, while offering features like async distributed checkpointing for building an end-to-end production workflow.
6 Conclusion
TorchTitan is a powerful and flexible framework for training LLMs. It offers composability, allowing users to combine various parallelism techniques (FSDP, TP, and PP), memory optimization methods (Float8 and activation checkpointing), and integration with torch.compile to optimize training efficiency. TorchTitan is highly flexible, adaptable to evolving model architectures and hardware advancements, and features a modular design with multi-axis metrics that foster innovation and experimentation. TorchTitan also prioritizes interpretability, production-grade training, and PyTorch native capabilities. Additionally, it provides high-performance training with elastic scalability, comprehensive training recipes and guidelines, and expert guidance on selecting and combining distributed training techniques. As shown in the experiment sections, TorchTitan provides training accelerations of 65.08% with 1D parallelism at the 128-GPU scale (Llama 3.1 8B), an additional 12.59% with 2D parallelism at the 256-GPU scale (Llama 3.1 70B), and an additional 30% with 3D parallelism at the 512-GPU scale (Llama 3.1 405B) over optimized baselines. With its robust features and high efficiency, TorchTitan is an ideal one-stop solution for challenging LLM training tasks.
7 Acknowledgements
We thank Soumith Chintala, Gregory Chanan, and Damien Sereni for their leadership support and product guidance. We thank Vasiliy Kuznetsov, Driss Guessous, Ke Wen, Yifu Wang, Xilun Wu, Liang Luo, and Gokul Gunasekaran for contributing fixes to the TorchTitan repository. Finally, we would like to thank our partners Linsong Chu and Davis Wertheimer at IBM Research for evaluating TorchTitan as a production platform and providing us with invaluable feedback.